OpenStudy (anonymous):

How to solve 2cos^4(x)-sin^2(x)=0? I don't know which identities to use

OpenStudy (anonymous):

You can simplify the left hand using \[\cos^2x+\sin^2x=1\] and treat it like a polynomial and go from there

OpenStudy (anonymous):

probably want to turn \(\cos^4(x)\) in to something with \(\sin^2(x)\) is my first guess

OpenStudy (anonymous):

oh maybe i am wrong

OpenStudy (anonymous):

would i move sin^2x to the other side in order to factor out a cos^2x?

jimthompson5910 (jim_thompson5910):

This is one way to do it \[\Large 2\cos^4(x)-\sin^2(x)=0\] \[\Large 2\cos^4(x)-(1-\cos^2(x))=0\] \[\Large 2\cos^4(x)-1+\cos^2(x)=0\] \[\Large 2\cos^4(x)+\cos^2(x)-1=0\] \[\Large 2(\cos^2(x))^2+\cos^2(x)-1=0\] \[\Large 2z^2+z-1=0 ... \text{Let } z = \cos^2(x)\] Use the quadratic formula to solve for z. Then use the solutions for z to find the solutions for x.

OpenStudy (anonymous):

thanks @jim_thompson5910 , i didn't look at it as a quadratic before! and thanks to everybody else too for helping

jimthompson5910 (jim_thompson5910):

you're welcome